A neonate weighs 2012 g.He is to recieve ampicillin 100 to 200 mg/kg/24 h divided q 6 h iv per syringe for otitis media. a. Childs weight is kg. b.what 24 hour dosage range is need? c.what single -dose range is needed?
Answer #1

ANSWER.

a) Weight of the child in kg.

         Weight of the child in gm = 2012 gm.

                 1 kg = 1000 mg.

     So

        Weight of the child in kg = 2012 / 1000

                                               = 2.012 kg.

                                              = 2 kg.

b) Calculation of 24 hr dose range.

           Ordered dose = 100 to 200 mg/kg/24 hr.

           Weight of the patient = 2 kg.

     So

        Dose required for a patient with 2 kg body weight = 100 * 2 to 200 * 2

                                                                               = 200mg to 400 mg / 24 hr.

c) Calculation of single dose range.

            Dose required for 24 hr = 200 mg to 400 mg.

         Dose for 24 hrs given as q6h (4times a day) as per order.

             So

               Single dose range = 200/4 to 400/4

                                             = 50 mg to 100 mg.